OpenStudy (anonymous):

solve A in the given equation cos^2A=1-COS^2A

OpenStudy (anonymous):

2cos 2A = 1 ==> cos 2A = 1/2 ==> 2A=(pi/6)+2(pi)k

OpenStudy (anonymous):

cosine square 2a equals to 1 minus cosine square 2a

OpenStudy (anonymous):

WHAT DOES pi means

OpenStudy (anonymous):

3.14...

OpenStudy (anonymous):

and k?

OpenStudy (anonymous):

so the answer is?

OpenStudy (anonymous):

k is an integer

OpenStudy (anonymous):

k\[k \in Z\]

OpenStudy (anonymous):

ahh whats the value if A?

OpenStudy (anonymous):

of a

OpenStudy (anonymous):

ahh thanks !! ^^

OpenStudy (anonymous):

\[sorry ( \pi/12) +\pi k\]

OpenStudy (anonymous):

i hope i helped

OpenStudy (anonymous):

my answer was 1-tan^x ?

OpenStudy (anonymous):

A is an angle